### v1

# Saner: Annotation-free Societal Attribute Neutralizer for Debiasing CLIP

Abstract
Large-scale vision-language models, such as CLIP, are known to contain harmful societal bias regarding protected attributes (e.g., gender and age). In this paper, we aim to address the problems of societal bias in CLIP. Although previous studies have proposed to debias societal bias through adversarial learning or test-time projecting, our comprehensive study of these works identifies two critical limitations: 1) loss of attribute information when it is explicitly disclosed in the input and 2) use of the attribute annotations during debiasing process. To mitigate societal bias in CLIP and overcome these limitations simultaneously, we introduce a simple-yet-effective debiasing method called SANER (societal attribute neutralizer) that eliminates attribute information from CLIP text features only of attribute-neutral descriptions. Experimental results show that SANER, which does not require attribute annotations and preserves original information for attribute-specific descriptions, demonstrates superior debiasing ability than the existing methods.
Keywords:
Societal bias CLIP Debiasing Fairness
1 Introduction
Figure 1: Our debiasing method, SANER, overcomes the limitations in existing methods: (a) attribute information is retained after debiasing, and (b) protected attribute annotations are not required for debiasing.
Large-scale vision-language models (VLMs), such as CLIP [37], have demonstrated a remarkable capability in multi-modal understanding [32, 50] and generation [39, 48, 57], being trained with million-scale image-text pairs. Utilizing these VLMs, recent vision models have achieved significant performance enhancements across a wide range of computer vision tasks (e.g., captioning [15, 33, 58, 28, 56] and object detection [27, 60]), without the necessity for task-specific training [45].
Despite the success, several works have identified societal bias regarding demographic attributes, such as gender and age, in these VLMs [54, 18, 36, 5, 21, 1, 55], potentially causing unfair or prejudicial decisions by models. Hall et al. [20] conducted audits on performance disparity, particularly with respect to gender, and revealed gender-dependency of the CLIP performance. Qiu et al. [36] also demonstrated that adopting CLIP for caption evaluation tends to favor gender-stereotypical sentences (e.g., preferring “A woman is cooking” over “A man is cooking” for images depicting men), highlighting the inherent gender bias. These findings underscore the importance of addressing bias in VLMs.
Some studies have proposed to mitigate societal bias in VLMs [4, 43, 11, 9]. Adversarial debiasing [4, 43, 11] fine-tunes CLIP to lessen leakage of protected attributes [^1] We refer to any demographic variables, like age and gender, as protected attribute (or attribute in short), based on which a model’s decisions should not be made. into the features, while projection-based debiasing [9] removes the protected attribute encoded in CLIP features in the inference phase. Our holistic review of these pioneering works (Sec. 3), though, identifies the following potential disadvantages or controversies in their design choices.
Loss of attribute information explicitly disclosed in the input. Some methods aim to completely remove attribute information by decorrelating the attribute and the features [11] or by squashing the subspace associated with the attribute [9], even when the attribute is explicitly disclosed. This choice can limit the generalizability of a VLM’s features to a spectrum of downstream tasks (e.g., Stable Diffusion [39] generates male images for text prompt “a female doctor” when encoded with debiased CLIP as shown in Fig. 1 (a) and Sec. 6.2), while it works for attribute-agnostic downstream tasks [26, 35].
Use of the attribute annotations. Adversarial debiasing methods [4, 43, 11] require protected attribute annotations, as provided in FairFace [24], for fine-tuning (Fig. 1 (b)). Datasets with attribute annotations are still scarce, perhaps partly because the annotation process needs ethical considerations [2], limiting their applicability and generalizability. This dataset scarcity also causes the limited diversity of images and text descriptions used for fine-tuning the VLM, potentially inducing overfitting.
This paper presents a simple-yet-effective debiasing approach for CLIP, called SANER (societal attribute neutralizer), that simultaneously overcomes the aforementioned limitations. Specifically, SANER trains a debiasing layer (i.e., a multilayer perception) to amend CLIP text feature vectors of attribute-neutral descriptions, given by attribute neutralization, such that they are equidistant to those of attribute-specific descriptions using annotation-free debiasing loss. With this, only feature vectors for attribute-neutral descriptions are debiased, whereas the attribute-specific ones retain the original information. Attribute-specific descriptions for all possible attribute groups [^2] Attribute group is a class in a protected attribute (e.g., female and male in gender). can be easily augmented by modifying the attribute-specific words in the original descriptions, directing the training without attribute annotations.
Contribution. Thanks to our annotation-free debiasing pipeline, SANER is designed to be compatible with any dataset of image-text pairs, such as COCO [29]. This provides denser guidance for training the debiasing layer compared to the existing methods. Experiments on both discriminative and generative tasks (i.e., text-to-image retrieval [19] and text-to-image generation [39]) show that SANER can mitigate gender, age, and racial biases of CLIP. Moreover, we demonstrate that SANER outperforms the existing methods [4, 9], showing that SANER leads to less attribute-dependency of the downstream performance while overcoming the limitations in existing methods.
2 Related Work
Societal bias in VLMs. With the growing application of VLMs, such as CLIP [37], across various tasks, there has been increasing awareness of their inherent social biases [20, 12, 51, 40, 55, 41, 46]. Hall et al. [20] examined gender bias in CLIP, uncovering significant discrepancies in object recognition performance based on the gender depicted in images. For example, the accuracy in predicting an umbrella when it appears with women is much better than with men.
The existence of such societal bias poses a substantial risk of perpetuating and exacerbating discrimination against historically marginalized groups [5, 36, 47, 53, 22]. Birhane et al. [5] demonstrated that adopting CLIP-based filtering in the dataset creation causes a risk of selecting stereotypical images regarding specific demographic subgroups. For instance, for an image of a female astronaut, the caption “This is a photograph of a smiling housewife” is favored over “This is a portrait of an astronaut” by CLIP, resulting in datasets that contain a number of image-caption pairs with harmful stereotypical descriptions.
Debiasing societal bias in CLIP. Recognizing the problem of bias in CLIP, some recent research has concentrated on mitigating societal bias in CLIP [4, 43, 11, 9]. These debiasing efforts intend to remove attribute information from CLIP features, resulting in less biased prediction outcomes. For example, Berg et al. [4] proposed a prompt tuning-based debiasing method based on an adversarial learning paradigm, minimizing the encoding of protected attribute information by the CLIP text encoder. However, in Sec. 3, we identify that these methods contain critical limitations that pose challenges in utilizing various datasets for debiasing and applying debiased models to a wide range of downstream tasks.
3 Review: Existing Debiasing Methods
Several debiasing approaches for CLIP have been introduced, broadly categorized into two main types: adversarial debiasing [4, 43, 11] and projection-based debiasing [9]. This section conducts an in-depth analysis of these existing debiasing strategies, highlighting their respective limitations.
Notation. Let $\mathcal{D}$ denote a dataset, each of whose sample is quadruple $(v,t,a,d)$, where $v$ is an image, $t$ is a text description, $a\in\mathcal{A}$ is a protected attribute annotation from set $\mathcal{A}$ of all attribute groups, and $d$ is the ground-truth annotation for a downstream task (if any). The CLIP text and image encoders, denoted by $f_{\text{t}}(t)\in\mathbb{R}^{K}$ and $f_{\text{v}}(v)\in\mathbb{R}^{K}$, respectively, take $t$ and $v$ as input and generate corresponding feature vectors in a common space.
3.1 Adversarial debiasing
Adversarial debiasing [4, 43, 11] aims to eliminate protected attribute information in the CLIP features. Specifically, an adversarial classifier is employed to predict and remove protected attribute $a$ from CLIP features.
[…]
$s_{c}(v)=f_{\text{v}}(v)^{\top}f_{\text{t}}(t_{c}^{\prime}).$ (1)
Let $s(v)\in\mathbb{R}^{|\mathcal{C}|}$ denote a vector, each of whose elements is the similarity score $s_{c}(v)$ for a concept in $\mathcal{C}$. Due to the attribute non-specificity of concepts in $\mathcal{C}$, $s(v)$ should not correlate with attribute $a$ of $v$. However, the CLIP text encoder can embed $a$ into $s(v)$ due to bias, allowing an attribute classifier to predict $a$. We denote the probability of being $a$ given $s(v)$ (or a prediction score of the attribute classifier) by $m_{a}(s(v))$. Prompt tuning-based debiasing [4] uses $m_{a}$ for adversarial loss, given by
$\mathcal{L}_{\text{adv}}=-\sum_{v\in\mathcal{D}}\log m_{a}(s(v)).$ (2)
Minimizing $\mathcal{L}_{\text{adv}}$ with respect to $l$ reduces attribute leakage through $s(v)$. A contrastive loss between image and text features is also used for regularization.
The experiments [4] showed that this method could effectively reduce attribute leakage through $f_{\text{t}}(t)$, but the limited number of concepts [^3] Their experiments used $10$ concepts. may limit downstream tasks that enjoy the debiased features because as $l$ is learned only through a sparse set of concepts. Additionally, attribute annotations are necessary for the adversarial loss, resulting in the exclusive use of face-centric image datasets (e.g., FairFace [24]) as $\mathcal{D}$.
Additive Residual Learner (ARL) [43] is designed to remove attribute information from CLIP image features. This method assumes that a debiasing layer [^4] A fully-connected layer is used. $r$ can identify a vector to additively amend attribute-neutral image feature vector $\delta(v)$, i.e.,
$\delta(v)=f_{\text{v}}(v)-r(f_{\text{v}}(v)).$ (3)
Similarly to [4], an adversarial classifier is trained to predict $a$ from $\delta$ with adversarial loss
$\mathcal{L}_{\text{adv}}=-\sum_{v\in\mathcal{D}}\log m_{a}(\delta(v)).$ (4)
The reconstruction loss between $f_{\text{v}}(v)$ and $\delta(v)$ regularizes the training to maintain the original features.
This method shares a common limitation with prompt tuning-based debiasing [4], notably requiring attribute annotations. Another limitation is that it tries to remove attribute features even when attributes of people in images are explicitly disclosed (i.e., when the person is depicted in an image). Consequently, debiased CLIP is ignorant of protected attributes.
Mapper [11] aims to reduce spurious correlations between protected attributes $a$ and task label $d$ in $\mathcal{D}$. It applies mappings $f_{\text{v}}^{\prime}$ and $f_{\text{t}}^{\prime}$ to image and text features, respectively, as $x_{\text{v}}(v)=f_{\text{v}}^{\prime}(f_{\text{v}}(v))$ and $x_{\text{t}}(t)=f_{\text{t}}^{\prime}(f_{\text{t}}(t))$ for mitigating dependence on $a$. The adversarial loss is computed using a dependence measure $\text{Dep}(\cdot,\cdot)$ to quantify statistical dependence between features as:
$\mathcal{L}_{\text{adv}}=-\text{Dep}(x_{\text{v}}(v),a)-\text{Dep}(x_{\text{t}}(t),a).$ (5)
These mapping functions are also trained to maximize the statistical dependence between the features after the mapping and task label $d$, i.e.,
$\text{Dep}(x_{\text{v}}(v),d)+\text{Dep}(x_{\text{t}}(t),d),$ (6)
to retain the predictive power on the downstream task while reducing bias.
Similar to [4, 43], Mapper relies on attribute annotations. Moreover, it is designed only to address the spurious correlations between the attribute and task labels for a specific task but not for different tasks.
3.2 Projection-based debiasing
[…]
$P=I-U(U^{\top}U)^{-1}U^{\top}$ (7)
where $I$ is the identity matrix. $P$ can project a CLIP text feature vector $f_{\text{t}}(t)$ for a text description $t$ into the orthogonal complement by $Pf_{\text{t}}(t)$.
Unlike adversarial debiasing, which requires training by gradient descent update, $P$ has a closed-form solution and is computed in the inference phase. However, as with ARL, this method also eliminates attribute information even from descriptions with explicit attributes (e.g., “A photo of a female doctor”).
3.3 Summary of the challenges
The existing debiasing methods, including adversarial and projection-based, reveal several challenges: 1) Loss of attribute information (ARL and projection-based) even with explicit attribute description narrows down the utility of the debiased CLIP. For example, a text-to-image generative model with gender-debiased CLIP features may not properly depict explicitly specified gender (as
shown in Sec. 6.2), 2) Dependency on attribute annotations (prompt tuning, ARL, and Mapper) constrains the range of datasets that can be utilized, often necessitating the use of face-centric image datasets (e.g., FairFace), as opposed to more diverse, natural image datasets (e.g., COCO [29]).
4 Societal Attribute Neutralizer (SANER)
Figure 2: An overview of SANER, exemplified by binary gender.
Our method for debiasing CLIP features, SANER, addresses the limitations of the existing methods identified in Section 3. Notably, SANER 1) retains attribute information in cases where the person’s attributes are explicitly described and 2) eliminates the reliance on attribute annotations, allowing the use of any image-text dataset for training the debiasing layer.
SANER comprises 1) attribute neutralization, which eliminates protected attribute information from input text (Section 4.1); 2) feature modification, which removes attribute information from the CLIP text features by amending them with a debiasing layer (Section 4.2); 3) attribute annotation-free debiasing loss, ensuring the features are not biased towards any attribute group $g\in\mathcal{A}$ (Section 4.3); and 4) regularization losses, which preserve the original CLIP features and the alignment between image and text features (Section 4.4).
Figure 2 shows an overview of SANER. We train the debiasing layer for feature modification over an arbitrary dataset $\mathcal{D}=\{(v,t)\}$ of image $v$ and text description $t$ (e.g., image caption, alt text) pairs, which does not provide attribute annotation $a$ as well as target task label $d$.
4.1 Attribute neutralization
We first modify text description $t\in\mathcal{D}$ that contain person-related words, [^5] The person-related words include any words relevant to a person (e.g., person, girl, man). The complete list is available in the supplementary material. to remove attribute-specific words. Taking binary gender as a protected attribute, [^6] Following prior research [4, 43, 11, 9, 59, 52, 6], we focus on the binary gender but recognize the importance of inclusivity. SANER applies to non-binary genders. i.e., $\mathcal{A}=\{\texttt{female},\texttt{male}\}$, as example, the text description
[…]
contains attribute information (i.e., woman). We replace attribute-specific terms [^7] We use gender words defined in [23]. with the attribute-neutral ones to obtain an attribute-neutral text:
$\xi_{\text{n}}(t)=\text{``A \text@underline{person} is eating salad.''}$.”
where $\xi_{\text{n}}$ denotes a function for attribute neutralization. Neutralization can be done for other attributes, such as age. [^8] Examples for the race attribute are in the supplementary material. We remove age-specific terms [^9] We define age-specific terms. The list is in the supplementary material. (e.g., young and senior) in text descriptions, for instance, “A young woman is eating salad” $\rightarrow$ “A woman is eating salad”. In contrast to the previous approach [9], which is optimized not to predict the attribute information from the original description $t$, we target the attribute-neutral descriptions $\xi_{\text{n}}(t)$ to preserve the attribute information in the features of attribute-specific descriptions.
4.2 Feature modification
CLIP text features $z(t)=f_{\text{t}}(\xi_{\text{n}}(t))$ after attribute neutralization can still convey the protected attribute information due to CLIP’s bias. To remove such bias, we append a learnable debiasing layer $r$ on top of $f_{\text{t}}$, inspired by recent CLIP fine-tuning techniques [43, 17]. Neutralized $t$’s debiased feature $h(t)$ is given by
$h(t)=z(t)+r(z(t)).$ (8)
4.3 Attribute annotation-free debiasing loss
To train $r$ to extract attribute information from CLIP features without attribute annotations, we generate a set $\mathcal{T}$ of attribute-specific descriptions for $t\in\mathcal{D}$ and for $g\in\mathcal{A}$, i.e., $\mathcal{T}=\{\xi_{g}(t)|t\in\mathcal{D},g\in\mathcal{A}\}$, where $\xi_{g}(t)$ generates a description specific to attribute group $g$ from $t$. For the binary gender case, this involves generating descriptions with female- and male-specific words. For instance, from the text description, “A woman is eating salad.”, we generate two sentences with female and male attributes:
A woman is eating salad.
A man is eating salad.
The debiasing loss trains $r$ such that $h(t)$ is equidistant from $f_{\text{t}}(\xi_{g}(t))$ for all attribute groups in $\mathcal{A}$, ensuring an impartial representation across the spectrum of attribute groups. We implement this loss as the standard deviation of the cosine similarity between $h(t)$ and $f_{\text{t}}(\xi_{g}(t))$. Let $s_{g}(t)$ denote the similarity, i.e.,
$s_{g}(t)=\frac{h(t)^{\top}f_{\text{t}}(\xi_{g}(t))}{\|h(t)\|\;\|f_{\text{t}}(\xi_{g}(t))\|}.$ (9)
The debiasing loss $\mathcal{L}_{\text{deb}}$ is defined as
$\mathcal{L}_{\text{deb}}=\sqrt{\frac{1}{|\mathcal{D}|}\sum_{t\in\mathcal{D}}(s_{g}(t)-\bar{s}(t))^{2}},$ (10)
where $\bar{s}(t)=\sum_{g\in\mathcal{A}}s_{g}(t)/|\mathcal{A}|$. A lower standard deviation means $s_{g}$ is close to $\bar{s}$, leading to $h(t)$ being equidistant to $f_{\text{t}}(\xi_{g}(t))$ for all $g\in\mathcal{A}$. Notably, this debiasing loss can be computed without attribute annotations.
4.4 Regularization losses
Applying the debiasing loss alone significantly changes original CLIP features, thereby losing semantics [4, 43]. To maintain the alignment of resulting image-text features, we utilize reconstruction loss [43] and contrastive loss [37]. Reconstruction loss $\mathcal{L}_{\text{recon}}$ is the mean squared error between $f_{\text{t}}(t)$ and $h(t)$. Contrastive loss $\mathcal{L}_{\text{cont}}$ aims to minimize the negative log-likelihood of input image-caption pairs, $f_{\text{v}}(v)$ and $f_{\text{t}}(t)$, in comparison to negative ones.
4.5 Training and inference
The overall loss $\mathcal{L}$ is given by:
$\mathcal{L}=\alpha\mathcal{L}_{\text{deb}}+\beta\mathcal{L}_{\text{recon}}+\gamma\mathcal{L}_{\text{cont}},$ (11)
where $\alpha$, $\beta$, and $\gamma$ are the hyperparameters to weight respective losses.
During inference, we apply the trained debiasing layer $r$ and use the modified text features $r(f_{\text{t}}(t))$ as the CLIP text features.
5 Experiments: Text-to-Image Retrieval
Following previous studies [4, 43, 11, 9], we evaluate SANER on the text-to-image retrieval task regarding gender, age, and racial biases. Further analysis, such as the ablation study of the loss components, is in the supplementary materials.
5.1 Experimental settings
Evaluation metric. We employ the MaxSkew metric [19], utilized in the previous studies [4, 43, 11, 9], to quantify the societal bias in CLIP in the text-to-image retrieval task. MaxSkew measures the disparity between the attribute distribution of $|\mathcal{A}|$ protected attributes in the top-$k$ retrieved images. Let $\eta_{ak}(q)$ denote the ratio of images labeled with attribute $a$ in the top-$k$ retrieved images. For attribute neutral query $q$, $\eta_{ak}(q)$ should be $1/|\mathcal{A}|$ if the model is unbiased. MaxSkew@$k$ is defined as:
[…]
Ideally, MaxSkew@$k$ is 0 but is larger when a model biased.
Evaluation setting. For the attribute-neutral queries, we use template-based queries such as “a photo of a $c$ person”, where $c$ is an attribute-neutral concept. Prior work [4] has defined a list of person-related adjectives, such as clever and attractive, as attribute-neutral concepts. We extend the list to encompass occupations (e.g., doctor, nurse), and activities (e.g., cooking and cleaning) for a more comprehensive evaluation. MaxSkew@$k$ is computed per concept. [^10] The complete lists of the concepts are in the supplementary material.
We also evaluate zero-shot image classification accuracy on ImageNet-1K [42] to ensure that debiasing does not spoil the original CLIP’s performance.
Evaluation datasets. We utilize two datasets, FairFace [24] and PATA [43], which consist of images alongside protected attribute annotations (e.g., female and male for gender) associated with the person in each image. FairFace consists of $10,954$ cropped face-centric images, while PATA contains $4,934$ natural images with a single person. Most debiasing approaches only report the performance on the FairFace dataset, but we additionally employ PATA to evaluate the debias performance on more diverse images.
Methods for comparison. We compare SANER against existing methods, i.e., prompt tuning-based debiasing [4] and projection-based debiasing [9], whose code is publicly available. Unfortunately, we could not reproduce the other methods[43, 11] since sufficient reproduction details are unavailable.
Implementation details. Following the previous works [4, 9], we employ CLIP [37] with ViT-B/16 backbone [14] as a target model in our experiments. We train the debiasing layer (a multilayer perception with two linear layers with ReLU activation [34]) appended to the model using $170,624$ image-caption pairs, which is a subset of the COCO training set [29] with person-related words/phrases (e.g., person and boy). We empirically set $\alpha$, $\beta$, and $\gamma$ to $1.0$, $0.1$, and $0.0001$ (Eq. 11), respectively, and train for $5$ epochs. Further details are provided in the supplementary materials.
Table 1: Gender bias, evaluated by MaxSkew@1000 (scaled by $100$), on FairFace and PATA for the original CLIP (Original), prompt tuning-based debiasing (Prompt), projection-based debiasing (Projection), and our method (SANER). A lower value is better (less gender bias). Bold represents the best across the models.
FairFace PATA
CLIP Model Adjective Occupation Activity Adjective Occupation Activity
Original [37] $22.9$ $33.7$ $19.5$ $12.1$ $18.7$ $10.7$
Prompt [4] $12.3$ $29.9$ $20.0$ $6.7$ $16.5$ $10.2$
Projection [9] $15.4$ $37.4$ $15.0$ $6.4$ $13.6$ $5.4$
SANER (Ours) $\mathbf{8.9}$ $\mathbf{14.5}$ $\mathbf{7.7}$ $\mathbf{5.4}$ $\mathbf{9.5}$ $\mathbf{3.3}$
Table 2: Age bias, evaluated by MaxSkew@1000 (scaled by $100$), on FairFace and PATA. Bold denotes the best across the models.
FairFace PATA
CLIP Model Adjective Occupation Activity Adjective Occupation Activity
Original[37] $111.1$ $121.1$ $113.0$ $40.4$ $44.4$ $39.7$
Projection [9] $107.6$ $112.8$ $\mathbf{100.0}$ $37.6$ $45.6$ $45.5$
SANER (Ours) $\mathbf{96.0}$ $\mathbf{112.6}$ $101.9$ $\mathbf{30.3}$ $\mathbf{36.7}$ $\mathbf{27.5}$
Table 3: Racial bias, evaluated by MaxSkew@1000 (scaled by $100$), on FairFace and PATA. Bold denotes the best across the models.
FairFace PATA
CLIP Model Adjective Occupation Activity Adjective Occupation Activity
Original[37] $62.2$ $57.4$ $68.3$ $33.4$ $28.6$ $31.5$
Projection [9] $56.9$ $75.3$ $67.0$ $\mathbf{19.9}$ $43.8$ $26.3$
SANER (Ours) $\mathbf{49.3}$ $\mathbf{45.7}$ $\mathbf{46.6}$ $28.9$ $\mathbf{21.2}$ $\mathbf{20.5}$
5.2 Gender bias analysis
Table 1 summarizes MaxSkew@1000 on FairFace and PATA for gender bias, showing that SANER mitigates gender bias the most among all methods. In contrast to the existing methods, this tendency is consistent across 1) the datasets with different image domains (i.e., face-centric and natural images) and 2) concept types (i.e., adjective, occupation, and activity). For instance, while the prompt tuning-based method [4] fails to mitigate bias for activity concepts on FairFace (i.e., bias is amplified from $19.5$ to $20.0$), SANER significantly mitigates bias from $19.5$ to $7.7$. This verifies the better debiasing performance of SANER compared to the existing methods on diverse concepts and image domains, possibly because SANER is trained with diverse text descriptions (i.e., captions in COCO), which are not constrained like pre-defined concepts required in the previous method [4].
Insight: Our debiasing method SANER, which does not necessitate attribute annotations, outperforms previous methods, including those that require annotations, in mitigating gender bias.
5.3 Age and racial biases analysis
Tables 2 and 3 present the results of MaxSkew@1000 for age and racial biases, respectively. We compare SANER with projection-based debiasing [9], as the prompt tuning-based method [4] does not provide age and race debiasing variants. Similarly to the results for gender bias, SANER surpasses the existing method across the datasets and the concept types. For example, SANER successfully mitigates the age bias on PATA. Meanwhile, the projection-based method works only on the adjective concepts and amplifies the bias on the other concepts (e.g., from $39.7$ to $45.5$). These results validate the generalizability of SANER in bias mitigation across the protected attributes.
Insight: SANER also demonstrates superior performance in mitigating age and racial biases compared to the existing method.
5.4 Zero-shot image classification
Table 4: Accuracy on Image-Net-1K.
CLIP Model Acc. (%)
Original [37] $65.4$
Prompt [4] $64.1$
SANER (Ours) $65.2$
To verify whether debiasing harms the zero-shot image classification performance of the original CLIP, we evaluate the prompt tuning-based method [4] and SANER on ImageNet-1K in terms of classification accuracy. Projection-based debiasing [9] does not apply to this evaluation because zero-shot prompts, such as “a photo of a car,” do not necessarily include person-related words. [^11] Projection-based debiasing requires modified input text to include attribute terms. The results are shown in Tab. 4, showing that applying SANER maintains classification performance, whereas the performance of the prompt tuning-based method slightly degrades.
Insight: SANER does not compromise the classification performance of the original CLIP.
6 Experiments: Text-to-Image Generation
We also evaluate SANER on text-to-image generation, for which societal bias is actively investigated [3, 7, 30, 31]. Specifically, we conduct two experiments from different aspects: 1) gender bias regarding occupations using gender-neutral prompts (Sec. 6.1), and 2) retention of attribute information when prompts explicitly disclose gender (Sec. 6.2).
Image generation settings. We use Stable Diffusion (SD) v2.1 [39] as the text-to-image generation model. The CLIP text encoder in SD is replaced with the debiased one for evaluation. Following [9], we use gender-neutral prompts with specifying occupations to analyze gender bias in generated images. These prompts are derived from the template “A photo of a $o$”, where $o$ is replaced with specific occupation terms, such as doctor and teacher. [^12] The list of the occupations is in the supplementary material. On the other hand, we employ gender-specific prompts to evaluate the capability of attribute information retention. Concretely, a gender term, either female or male, is added just before the occupation terms, i.e., “A photo of a {female/male} $o$”, to see if generated images specify the gender. We generate $100$ images for each prompt with SD’s default hyperparameters shown in [39].
Evaluation metrics. For the bias evaluation for the generative task, we use statistical parity (SP) metric that measures the disparity of attribute groups in generated images [8, 49, 9]. Specifically, we annotate binary gender labels (i.e., $\mathcal{A}=\{\texttt{female},\texttt{male}\}$) for the generated images with the assistance of human workers. [^13] Different from previous works that use pre-trained gender classifiers to assign gender labels, we do not use them due to their bias issues [38, 10, 13]. SP is defined as the difference between the empirical distribution $\kappa_{a}$ of gender $a$ and uniform distribution, given by:
$\text{SP}=\sqrt{\sum_{a\in\mathcal{A}}(\kappa_{a}-1/|\mathcal{A}|)^{2}},$ (13)
where $\kappa_{a}=N_{a}/\sum_{a^{\prime}}N_{a^{\prime}}$ with $N_{a}$ being the number of images annotated as $a$. For an unbiased text-to-image generation model, SP should be 0 but increases for biased models.
For gender-specific prompts for evaluating gender information retention, we compute the accuracy, i.e., the ratio of the images that contain the same gender as the prompt to all generated images.
6.1 Gender bias analysis
Table 5 summarizes SP scores, which demonstrate that applying SANER to Stable Diffusion notably mitigates gender bias regarding occupations. SANER again outperforms projection-based debiasing one (i.e., $0.39$ for SANER and $0.47$ for projection-based), highlighting the superiority of SANER.
We show visual examples in which SANER successfully mitigates gender bias in Fig. 3. For the gender-neutral prompt, “A photo of a designer,” the original SD and projection-based debiasing (Projection) predominantly generate images depicting a man. In contrast, ours shows a more balanced gender distribution.
Insight: Applying SANER debiased CLIP to Stable Diffusion results in generating images with a more gender-equal distribution.
6.2 Assessment of retention of attribute information
Table 5 also shows the accuracy regarding how much the gender of the person in generated images matches the gender specified in the prompts. For both prompts that describe women and men, using debiased CLIP by projection-based debiasing leads to losing gender information (i.e., accuracies for projection are much lower than those for the original). Conversely, SANER retains gender information (i.e., accuracies for SANER are $1.00$). Figure 4 confirms this, showing that using projection-based debiased CLIP results in generating male images for the prompt, “A photo of a female doctor”.
Insight: SANER retains gender information for input texts that explicitly include gender terms.
Table 5: Gender bias (SP) and retention of gender information (Accuracy) in images generated by Stable Diffusion with original CLIP text encoder (Original) and the model using projection-based debiased CLIP (Projection) and our debiased CLIP (SANER). We report the mean and standard deviation over the set of occupations. Female and male denote prompts with female and male specifications, respectively. Bold represents the best across the models.
Accuracy $\uparrow$
CLIP Model SP $\downarrow$ Female Male
Original [37] $0.51$$\pm 0.22$ $\mathbf{1.00}$$\pm 0.00$ $\mathbf{1.00}$$\pm 0.00$
Projection [9] $0.47$$\pm 0.19$ $0.58$$\pm 0.28$ $0.79$$\pm 0.18$
SANER (Ours) $\mathbf{0.39}$$\pm 0.22$ $\mathbf{1.00}$$\pm 0.00$ $\mathbf{1.00}$$\pm 0.00$
Figure 3: Generated images for the prompt, “A photo of a designer,” by the original Stable Diffusion (SD), projection-based debiased CLIP (Projection), and our debiased CLIP (SANER). We randomly sample $10$ images from generated images. Images framed in green denote those of the minority gender in the generated images (i.e., female).
Figure 4: Generated images for the prompt, “A photo of a female doctor,” by the original Stable Diffusion (SD), projection-based debiased CLIP (Projection), and our debiased CLIP (SANER). We randomly sample $10$ images from generated images. Red frame indicates images with incorrect gender (i.e., male).
7 Limitations
Bias metrics. While SANER outperforms the existing methods in mitigating gender, age, and racial biases according to existing bias metrics, such as MaxSkew@$k$ and SP, this does not guarantee that SANER is always superior to them in every aspect or scenario. As suggested in prior work [4, 43, 11, 9], reliance on these metrics provides a standardized measure of societal bias; however, there could be dimensions of bias not adequately captured by them, which implies the necessity of more holistic metrics to assess societal bias.
Further bias mitigation. The experimental results show that SANER noticeably mitigates societal bias in CLIP. Nonetheless, the bias is not completely eliminated (e.g., MaxSkew is not zero). A promising direction for further debiasing could involve debiasing the image encoder, specifically training a debiasing layer to remove attribute information from the visual features for images in which humans do not appear.
Intersectional bias analysis. Although we focus on gender, age, and racial biases separately in our experiments, following the previous debias works [4, 43, 11, 9], SANER can be easily extended to various protected attributes, including their combinations. For instance, considering the intersection of binary gender and age, we generate four sentences with (female, young), (female, old), (male, young), and (male, old) for the debiasing loss, e.g., “A young woman is eating salad” for the input text “A woman is eating salad”. This potential for addressing complex biases is noted in future research.
8 Conclusion
This paper proposed SANER, a simple-yet-effective debiasing method for CLIP, consisting of attribute neutralization and anotation-free debiasing loss. Consequently, SANER can use an arbitrary dataset of image-text pairs for training the debiasing layer, resulting in superior debiasing performance compared to the existing methods for both discriminative and generative tasks (i.e., text-to-image retrieval and text-to-image generation). We also confirmed that SANER retains attribute information for attribute-specific descriptions through the gender-specified prompts for text-to-image generation.
Future work can explore extending SANER to the vision encoder, removing attribute information from visual features in images without humans, and addressing intersectional biases, combining attributes like gender and age. Furthermore, SANER is applicable to any dataset of image-caption pairs, thus, scaling to larger datasets (e.g., Conceptual Captions [44]) for denser supervision is an interesting future direction.
Supplementary Material
This supplementary material includes:
  • Implementation details for SANER (Appendix 0.A).
  • Further analysis (Appendix 0.B).
  • List of person-, gender-, age-, and race-specific terms (Appendix 0.C).
  • List of concepts and occupations (Appendix 0.D).
  • Additional visual examples for image generation experiments (Appendix 0.E).
  • Potential negative impact (Appendix 0.F).
Appendix 0.A Implementation Details for SANER
We train the debiasing layer (a multilayer perception with two linear layers with ReLU activation) using $170,624$ image-caption pairs, which is a subset of the COCO training set with person-related words/phrases defined in Appendix 0.C. The hidden embedding dimensionality for the debiasing layer is set to $128$. We empirically set $\alpha$, $\beta$, and $\gamma$ to $1.0$, $0.1$, and $0.0001$ (Eq. (11) in the main paper), respectively. We set the training epochs, batch size, and learning rate to $5$, $128$, and $5\times 10^{-6}$, respectively.
Racial bias mitigation. For race attribute, we remove race-specific terms [^14] We define race-specific terms. The list is in Appendix 0.C. (e.g., African and Asian) in text descriptions, for instance, “An African woman is eating salad” $\rightarrow$ “A woman is eating salad”.
Appendix 0.B Additional Experiments
0.B.1 Loss ablation
To validate the effectiveness of each regularization loss (i.e., reconstruction loss and contrastive loss), we conduct an ablation study by removing one of the losses or both losses. Table 6 presents the results of gender bias. The results show that using both reconstruction and contrastive losses yields the best results regarding gender bias mitigation and zero-shot classification accuracy on ImangeNet [42]. Furthermore, SANER without regularization losses significantly degrades CLIP’s zero-shot classification ability (i.e., from $65.4$ to $58.8$). These observations confirm the importance of having both reconstruction and contrastive losses for the regularization.
Table 6: Gender bias, evaluated by MaxSkew@1000 (scaled by $100$), on FairFace and PATA for our method (SANER) with different regularization loss combinations. Recon denotes the use of the reconstruction loss, and cont represents the use of the contrastive loss. IN acc is the zero-shot classification accuracy on ImageNet. Adj, Occ, and Act represent the types of concepts (i.e., Adjective, Occupations, and Activity, respectively). A lower value is better (less gender bias). Bold represents the best across the SANER variants.
FairFace PATA
CLIP Model Adj Occ Act Adj Occ Act IN acc (%)
Original [37] $22.9$ $33.7$ $19.5$ $12.1$ $18.7$ $10.7$ $65.4$
SANER
(no regularization) $15.7$ $15.0$ $15.3$ $15.8$ $14.4$ $15.4$ $58.8$
Recon $10.2$ $16.7$ $11.4$ $6.2$ $12.0$ $6.2$ $64.2$
Cont $18.0$ $19.4$ $21.7$ $7.0$ $\mathbf{9.1}$ $9.0$ $63.0$
Recon+Cont $\mathbf{8.9}$ $\mathbf{14.5}$ $\mathbf{7.7}$ $\mathbf{5.4}$ $9.5$ $\mathbf{3.3}$ $\mathbf{65.2}$
0.B.2 Analysis on the data size
The experiments in the main paper (Sections 5 and 6) verify that SANER outperforms existing debiasing methods, showing a better bias mitigation ability in terms of gender and age biases. This superior performance of SANER may be because SANER is trained with diverse text descriptions (i.e., captions in COCO [29]), which are not constrained like pre-defined concepts required in the previous method [4]. In this section, we conduct an experiment to verify this hypothesis. Specifically, we use $n$ percent of the training samples (i.e., $17,624$ image-caption pairs of COCO) to evaluate the impact of the training dataset size. We use the same settings in Appendix 0.A, but use the different training epochs to align the number of iterations. The results are shown in Tab. 7.
The results validate that as the number of data samples increases, gender bias is reduced. Specifically, while using a part of the training data results in mitigating gender bias (i.e., MaxSkew scores are smaller than the original CLIP), using the full training samples (i.e., COCO-100%) gives the best results, showing the importance of the use of more diverse data for debiasing.
Table 7: Gender bias, evaluated by MaxSkew@1000 (scaled by $100$), on FairFace and PATA for the original CLIP (Original) and our method (SANER) with different data sizes. COCO-$n$denotes that we use $n$of the training samples. IN acc is the zero-shot classification accuracy on ImageNet. A lower value is better (less gender bias). Bold represents the best across the SANER variants.
FairFace PATA
CLIP Model Adj Occ Act Adj Occ Act IN acc (%)
Original [37] $22.9$ $33.7$ $19.5$ $12.1$ $18.7$ $10.7$ $65.4$
SANER
COCO-50% $15.7$ $19.4$ $17.9$ $10.0$ $14.1$ $10.2$ $\mathbf{65.4}$
COCO-75% $12.8$ $17.6$ $16.1$ $7.8$ $13.0$ $8.3$ $\mathbf{65.4}$
COCO-100% $\mathbf{8.9}$ $\mathbf{14.5}$ $\mathbf{7.7}$ $\mathbf{5.4}$ $\mathbf{9.5}$ $\mathbf{3.3}$ $65.2$
Appendix 0.C List of Person-, Gender-, Age-, Race-Specific Terms
The person-related words that are used to identify text descriptions that are relevant to humans (in Sec. 4.1 in the main paper) are as below:
actor, actress, adult, architect, artist, associate, aunt, baby, boy, boyfriend, brother, chairman, chairperson, chairwoman, chef, child, coach, colleague, comedian, counselor, cowboy, cowgirl, dancer, daughter, designer, director, doctor, driver, dude, elder, emperor, employee, employer, engineer, entrepreneur, executive, expecting, father, female, friend, gentleman, girl, girlfriend, guy, hairdresser, he, her, hers, herself, him, himself, his, husband, individual, infant, instructor, kid, lady, lawyer, leader, lecturer, male, man, manager, mechanic, member, mentor, mother, musician, neighbor, novelist, nurse, parent, partner, people, performer, person, pharmacist, photographer, physician, pilot, player, police officer, policeman, policewoman, politician, pregnant, prince, princess, professor, queen, relative, researcher, royal, scholar, scientist, secretary, server, she, sibling, singer, sister, son, specialist, spouse, student, surfer, surgeon, tailor, teacher, technician, teenager, their, theirs, them, themselves, therapist, they, toddler, uncle, veterinarian, volunteer, waiter, waitress, wife, woman, worker, writer, youth, and their plurals.
We list the gender-specific terms that are used to create attribute-neutral text descriptions $\xi_{n}(t)$ (in Sec. 4.1 in the main paper): woman, female, lady, mother, girl, aunt, wife, actress, princess, waitress, sister, queen, pregnant, daughter, she, her, hers, herself, man, male, father, gentleman, boy, uncle, husband, actor, prince, waiter, son, brother, guy, emperor, dude, cowboy, he, his, him, himself and their plurals (orange denotes female-specific words, and olive represents male-specific terms). To synthesize attribute-specific descriptions (i.e., $\mathcal{T}=\{\xi_{g}(t)|t\in\mathcal{D},g\in\mathcal{A}\}$ in Sec. 4.3 in the main paper) for binary gender, we replace person-specific terms in the attribute-neutral descriptions with their corresponding gender terms (e.g., person $\rightarrow$ woman and person $\rightarrow$ man).
We also list the age-specific terms used to create attribute-neutral text descriptions $\xi_{n}(t)$ (in Sec. 4.1 in the main paper): elderly ,baby, child, kid, teenager, adult, youth, infant, toddler, elder, girl, boy, young, old, teenage, and their plurals. To create attribute-specific descriptions (i.e., $\mathcal{T}=\{\xi_{g}(t)|t\in\mathcal{D},g\in\mathcal{A}\}$ in Sec. 4.3 in the main paper) for binary age, we add young or old just before the person-specific terms (e.g., person $\rightarrow$ young person and person $\rightarrow$ old person).
[…]
Appendix 0.D List of concepts and occupations
We show the concepts used to compute MaxSkew@$k$ (in Tab. 1 and 2 in the main paper). The list of the concepts for each category (i.e., adjectives, occupations, and activities) are as follows:
Adjectives: kind, successful, noncriminal, nonviolent, moral, rich, honest, smart, evil, unkind, unsuccessful, criminal, violent, immoral, poor, dishonest, dumb, clever, stupid, hardworking, lazy, nasty, trustworthy, untrustworthy, caring, dangerous, heroic, villainous, good. We refer these concepts from [4]. Regardig the templates used to create prompts, we use “This is the face of a {} person”, “This is a photo of a {} person”, “This person is {}”, “A {} person”, “A {} individual”, “This individual is {}”, “Photo of a {}”, “{}”, “This is a {}”, “This is the photo of a {}”, “A {}”, following [4]. The curly brackets are replaced with each concept.
Occupations: doctor, aerospace engineer, computer programmer, electrical engineer, scientist, artist, author, designer, musician, painter, photographer, singer, writer, architect, civil engineer, engineer, programmer, software developer, childcare worker, coach, dental assistant, dental hygienist, dentist, clerk, housekeeper, maid, massage therapist, nurse, therapist,psychologist, social assistant, social worker, teacher, professor, CEO, skateboarder, surfer, baseball player, football player, soccer player, tennis player. For the occupation list, we refer to [16] with some modifications such as additional occupations for more inclusive list. As for the prompt templates, we select ones that are suitable to occupations (e.g., “Photo of a {}”). We use this list for the text-to-image experiments with some modifications. Specifically, we remove the similar occupations (e.g., remove civil engineer as there is engineer). Additionally, we remove occupations where the text-to-image model [39] can not generate images with humans or it tends to generate multiple individuals.
Activities: arranging flowers, playing tennis, playing skateboarding, playing baseball, playing soccer, playing football, playing snowboarding, playing skiing, cleaning, dressmaking, tying tie, smiling, crying, laughing, cooking, making pizza, dancing, drinking beer, drinking wine, eating hotdog, eating cake, using computer, playing game, gardening, singing, petting dog, petting cat, makeup, shopping, playing piano, playing guitar, carrying baby. For activities, we use a subset of the Kinetics dataset [25] with some additional activities. For the prompt templates, we use “This is the face of a person who likes {}”, “ This is a photo of a person who likes {}”, “This person likes {}”, “A person who likes {}”, “Photo of a person who likes {}”, “This is a person who likes {}”.
Appendix 0.E Additional visual examples for image generation experiments
We show additional visual examples for Figures 3 and 4 in the main paper in Figures 5 and 6, respectively.
Figure 5: Generated images for the prompt, “A photo of a teacher,” by the original Stable Diffusion (SD), projection-based debiased CLIP (Projection), and our debiased CLIP (SANER). We randomly sample $10$ images from generated images. Images framed in green denote those of the minority gender in the generated images (i.e., male).
Figure 6: Generated images for the prompt, “A photo of a male musician,” by the original Stable Diffusion (SD), projection-based debiased CLIP (Projection), and our debiased CLIP (SANER). We randomly sample $10$ images from generated images. Red frame indicates images with incorrect gender (i.e., female).
Appendix 0.F Potential Negative Impact
Applying SANER to debias CLIP may lead to a potential negative impact where users might overlook remaining biases, assuming the process to be fully effective. While SANER performs better in gender and age bias mitigation, as evidenced by the MaxSkew and statistical parity metrics, this does not ensure that SANER mitigates all possible societal biases, and there could be dimensions of bias not adequately measured by these metrics. It is crucial to acknowledge that SANER, though impactful, is not an all-encompassing solution for removing societal bias. We notice that researchers must exercise due diligence in evaluating the application of SANER to avoid inadvertently introducing unanticipated biases.